# ChangeMamba: Remote Sensing Change Detection with Spatio-Temporal State Space Model

### v1

@@ V Conclusion
In this paper, we explore the potential of the emerging Mamba architecture for remote sensing image change detection tasks. Three architectures, MambaBCD, MambaSCD, and MambaBDA are developed for binary change detection, semantic change detection, and building damage assessment tasks, respectively. The encoders of all three architectures are cutting-edge visual Mamba, which can adequately learn the global contextual information of the input image with linear complexity. Then, for the decoder, in order to adequately learn about spatio-temporal relations, we propose three ways of modeling spatio-temporal relations which can adequately exploiting the attributes and advantages of Mamba architecture. Experiments on five benchmark datasets have fully revealed the potential of Mamba for multi-temporal remote sensing image data processing tasks. Compared to the CNN-based and Transformer-based approaches, the proposed three architectures can achieve SOTA performance on their respective tasks. Compared to some commonly used backbone networks, the Mamba architecture can extract features that are more suitable for downstream CD tasks. The combination of our proposed three spatio-temporal relationship modeling mechanisms and the Mamba architecture provides a new perspective on modeling spatio-temporal relationships, showing much better detection results on all three subtasks.

### v2

@@ V Conclusion
In this paper, we explore the potential of the emerging Mamba architecture for remote sensing image change detection tasks. Three architectures, MambaBCD, MambaSCD, and MambaBDA are developed for binary change detection, semantic change detection, and building damage assessment tasks, respectively. The encoders of all three architectures are cutting-edge Visual Mamba architecture, which can adequately learn the global contextual information of the input image with linear complexity. Then, for the change decoder, in order to adequately learn spatio-temporal features, we propose three ways of modeling spatio-temporal relationships which can adequately exploiting the attributes and advantages of Mamba architecture. Experiments on five benchmark datasets have fully revealed the potential of Mamba for multi-temporal remote sensing image data processing tasks. Compared to the CNN-based and Transformer-based approaches, the proposed three architectures can achieve SOTA performance on their respective tasks. Compared to some commonly used backbone networks, the Mamba architecture can extract features that are more suitable for downstream CD tasks. The combination of our proposed three spatio-temporal relationship modeling mechanisms and the Mamba architecture provides a new perspective on modeling spatio-temporal relationships, showing much better detection results on all three subtasks. Additional experiments show that our Mamba-based architecture is robust to degraded input data.